If 1.7 g of sucrose (C12H22O11) is dissolved in 250.0 mL of ethanol (density = 0.789 g/mL), what is the molality of the resulting solution?
1
Expert's answer
2019-09-16T04:33:00-0400

n(C12H22O11) = m(C12H22O11)/M(C12H22O11) = 1.7g / 342g/mol = 0.00497 mol.

m(C2H5OH) = V(C2H5OH) × ρ(C2H5OH) = 250.0 mL × 0.789 g/mL = 197.25 g.


0.00497 mol – 197.25 g

X mol          - 1000 g


X = 0.0252 mol ≈ 0,025 mol.

Therefore, the molality of the resulting solution is b = 0,025 mol/kg.